OpenStudy (adamaero):

MEDAL - Deposits x, at 3% compounded daily. How much interest after two years? Pe^rt ? xe^(.03[365]2) ? Is the two years computed correctly?

OpenStudy (adamaero):

@wahahaha

OpenStudy (anonymous):

hmmm I was thinking of doing it like this The EAR of computing it daily is either \[EAR=\left (1+\frac{.03}{365} \right )^{365}-1\] Or \(e^{.03}=3.045\) and then we square it for the 2 year period So it would be \[(e^{.03})^2\] or \[ \left(\left (1+\frac{.03}{365} \right )^{365}-1\right) ^2\]

OpenStudy (anonymous):

We are only interested in finding out the amount of overall interest Not what the amount deposited will equal in 2 years time

OpenStudy (adamaero):

It's with e for cont compounding. And, thanks, only interest...

OpenStudy (anonymous):

Ya so in this case you can use either cuz in reality the EAR for 365 pretty much equals e^r

OpenStudy (adamaero):

e^0.03 = 1.03 your EAR equared = 0.0009274

OpenStudy (anonymous):

hmmm u r wrong btw

OpenStudy (anonymous):

\[EAR=\left (1+\frac{.03}{365} \right )^{365}-1=.03045\] \[e^{r}-1=e^{.03}-1=.03045\]

OpenStudy (adamaero):

Oooooooo - 1